OpenStudy (anonymous):

Solve x2 + 10x = -21

OpenStudy (anonymous):

x = 7 and x = 3 x = -7 and x = -3 x = -7 and x = 3 x = 7 and x = -3

OpenStudy (anonymous):

x=-7,-3

OpenStudy (anonymous):

split 21 in such a way on addition or subtraction u get 10

OpenStudy (anonymous):

how did you get the answer

OpenStudy (anonymous):

x^2 +10x = -21 x^2+10x+21 = 0 x^2+7x+3x+21 =0 i assume u solve from here